# The Born approximation for the fixed energy Calderón problem

### v1

@@ 4.3. From the $A_{\kappa}$-amplitude to the Born approximation
    $q^{s}_{\kappa}-q\in\mathcal{C}(\overline{\mathbb{B}^{d}}\setminus\{0\})$ and $q^{s}_{\kappa}-q=0$ in $\partial\mathbb{B}^{d}$. If $q\in\mathcal{C}^{m}(\mathbb{B}^{d}\setminus\{0\})$ for some $m\in\mathbb{N}_{0}$ then $q^{s}_{\kappa}-q$ belongs to $\mathcal{C}^{m+1}(\mathbb{B}^{d}\setminus\{0\})$.
@@
(4.21) $\int_{\mathbb{R}_{+}}|A_{Q,\kappa}(t)|e^{-2sx}\,\mathrm{d}t=\frac{1}{\left\lvert\mathbb{S}^{d-1}\right\rvert}\int_{\mathbb{B}^{d}}|q^{s}_{\kappa}(x)||x|^{2(s-\nu_{d})}\,\mathrm{d}x.$
@@
$\lambda_{\ell}[q,\kappa]-\lambda_{\ell}[0,\kappa]=-m_{Q}\left(-(\ell+\nu_{d})^{2}\right)-m_{Q_{\kappa}}\left(-(\ell+\nu_{d})^{2}\right)\\ =-\int_{0}^{\infty}A_{Q,\kappa}(t)u_{\ell+\nu_{d}}(t,\kappa)^{2}\,\mathrm{d}t=-\int_{0}^{\infty}A_{Q,\kappa}(t)e^{-2\nu_{d}t}\frac{\varphi_{\ell}(\sqrt{\kappa}e^{-t})^{2}}{\varphi_{\ell}(\sqrt{\kappa})^{2}}\,\mathrm{d}t.$
@@
which proves (4.19) when $\kappa\notin\operatorname{Spec}_{H^{1}_{0}}(-\Delta)$. To see that the result holds only under the assumption that $\kappa$ is not a Dirichlet eigenvalue of $-\Delta+q$ notice that both sides of the above equations can be extended continuously to $\operatorname{Spec}_{H^{1}_{0}}(-\Delta)$ by (1.9) and Lemma 4.2.
Property (iii) follows from Theorem 4.4 (i) and (iii): $A_{Q,\kappa}-Q$ is continuous on $\mathbb{R}_{+}$ and can be extended continuously by zero to $[0,\infty)$. Then $q^{s}_{\kappa}-q$ is continuous on $\mathbb{B}^{d}\setminus\{0\}$, and can be continuously extended by zero to $\overline{\mathbb{B}^{d}}\setminus\{0\}$.
Part (iv) follows from estimate (4.15) by (4.17) and a change of variables. To see this, recall that $Q\in L^{1}(\mathbb{R}_{+})$ when $p>d/2$ (see Lemma 4.1), then $\alpha_{Q}(t)\leq\left\lVert Q\right\rVert_{L^{1}(\mathbb{R}_{+})}$ for all $t\in\mathbb{R}$. Hence, there exists $\beta_{q,\kappa}>0$ such that $|x|^{(2\alpha_{Q}(-\log|x|)+|\kappa|)}\leq|x|^{-\beta_{q,\kappa}}$.
@@ 5.1. Proof of 2 and 3
Therefore, by 1 we must have that
@@
which implies that $q^{\mathrm{B}}_{\kappa}\in\mathcal{B}_{d}$ (here we are using the identification of elements of $\mathcal{B}_{d}$ extended by zero, with elements of $\mathcal{E}^{\prime}_{\mathrm{rad}}(\mathbb{R}^{d})$). Finally, again by 1, $q^{\mathrm{B}}_{\kappa}$ satisfies (1.16). Note that once the existence of the Born approximation at energy $\kappa$ has been established, 3 follows from Proposition 4.6 and (5.3).
The proof of 2 will be concluded once we prove identity (1.18). The conditions on $\kappa$ and $q$ in the statement guarantee that the operator $\Lambda_{\kappa,q}-\Lambda_{\kappa,0}$ is well defined. Recall that $\mathcal{P}_{\ell,d}$ denotes the orthogonal projection of $L^{2}(\mathbb{S}^{d-1})$ onto $\mathfrak{H}_{\ell}^{d}$, and that $\mathcal{P}_{\ell,d}\overline{f}=\overline{\mathcal{P}_{\ell,d}f}$. Therefore we have that
@@
5.2. Proof of 5
By (4.18), (4.17) and (5.3) we have
@@ 5.2. Proof of 5
For the stability notice that if $\max_{j=1,2}\left\lVert q_{j}\right\rVert_{L^{p}(\{b<|x|<1\})}<K$, then one can choose a constant $M>0$ (dependent on $\kappa$) such that (4.16). Applying Theorem 4.5(ii), and letting $a=-\log b$, we obtain
@@ 5.3. Proof of 4
where $\mathcal{R}^{\mathbb{R}^{d}}_{0}(-z^{2})$ is the inverse of $-\Delta+z^{2}$ defined in $L^{2}(\mathbb{R}^{d})$. This operator in the whole space has the Green function
@@
$\left\lvert\left\langle q\overline{e_{\zeta_{1}}},\mathcal{R}_{q}(\kappa)(qe_{\zeta_{2}})\right\rangle_{L^{2}(\mathbb{B}^{d})}\right\rvert$ $\leq\left\langle\left\lvert q\overline{e_{\zeta_{1}}}\right\rvert,\left\lvert\mathcal{R}_{q}(\kappa)(qe_{\zeta_{2}})\right\rvert\right\rangle_{L^{2}(\mathbb{B}^{d})}$
@@
×$\quad\times\left(-\left(\frac{\left\lvert\operatorname{\mathrm{Re}}\zeta_{2}\right\rvert}{z}\right)^{\nu_{d}}+2\left\lvert\operatorname{\mathrm{Re}}\zeta_{2}\right\rvert K_{\nu_{d}}(2z)I_{\nu_{d}+1}(2\left\lvert\operatorname{\mathrm{Re}}\zeta_{2}\right\rvert)+2zK_{\nu_{d}+1}(2z)I_{\nu_{d}}(2\left\lvert\operatorname{\mathrm{Re}}\zeta_{2}\right\rvert)\right).$
@@
where $A_{Q}-A_{Q_{\kappa}}$ belongs to $L^{1}_{s}(\mathbb{R}_{+})$. Let $\psi_{z}(\cdot;\kappa)\in L^{2}(\mathbb{R}_{+})$ denote the Jost solution of (A.1). This solution is characterized by identities (A.2) and (A.3). Using Lemma A.3(i) we can write
@@
On the other hand, by Lemma A.2 and Lemma A.3, $\mathcal{G}_{\kappa}^{-1}-\operatorname{\mathrm{Id}}$ and $\mathcal{T}_{\kappa}-\operatorname{\mathrm{Id}}$ are both Volterra operators with a smooth kernel on $D$, and thus satisfy the conditions of Lemma A.1. Therefore $A_{Q,\kappa}-(A_{Q}-A_{Q_{\kappa}})$ belongs to $\mathcal{C}(\mathbb{R}_{+})$ always and to $\mathcal{C}^{m+1}(\mathbb{R}_{+})$ if $Q\in\mathcal{C}^{m}(\mathbb{R}_{+})$ with $m\in\mathbb{N}_{0}$; the analogous result for $A_{Q,\kappa}-Q$ follows immediately (recall that $Q_{\kappa}$, and hence $A_{Q_{\kappa}}$, are smooth). The fact that $A_{Q,\kappa}-Q$ can be extended continuously by zero to $[0,\infty)$ follows from estimate (4.15) that we now prove.
@@
Estimate (4.15) follows from (6.3) using the estimate
(6.4) $\left|A_{Q_{1}}(t)-Q_{1}(t)-\left(A_{Q_{2}}(t)-Q_{2}(t)\right)\right|\leq e^{2t(\alpha_{Q_{1}}(t)+\alpha_{Q_{2}}(t))}\int_{0}^{t}|Q_{1}(s)-Q_{2}(s)|\,\mathrm{d}s,$
which can be found in [DMMN24, Section 4] (recall that $\alpha_{Q}$ was defined in (4.14)).
It remains to prove property (iv). If $Q$ is positive the operator $H_{Q}=(-\partial_{t}^{2}+Q(t))$ has no negative eigenvalues nor a zero resonances as remarked in [RS00]. Then it follows from [RS00, Theorem 3] that $A_{Q}\in L^{1}(\mathbb{R}_{+})$. Also, we have $A_{Q_{\kappa}}\in L^{1}(\mathbb{R}_{+})$ if $\kappa<(j_{0,1})^{2}$ by Lemma 4.3, since $z_{\kappa}$ can be taken to be 00. It follows that $A_{Q,\kappa}\in L^{1}(\mathbb{R}_{+})$ by (6.1), since $\mathcal{G}_{\kappa}^{-1}\mathcal{T}_{\kappa}$ is bounded on $L^{1}(\mathbb{R}_{+})$ by Lemmas A.2 and A.3. This finishes the proof of the Theorem. ∎
@@
The first statement was proved by Simon in [Sim99, Theorem 1.5] for the case $\kappa=0$. The case $\kappa\neq 0$ follows from this and (6.1) by the local injectivity of the operators $\mathcal{G}_{\kappa}^{-1}\mathcal{T}_{\kappa}$, which satisfy that $\mathcal{G}_{\kappa}^{-1}\mathcal{T}_{\kappa}F|_{(0,a)}$ vanishes iff $F_{(0,a)}$ vanishes due to (6.2).
We now prove the stability estimate. Under the assumptions for $Q_{j}$ $j=1,2$ in the statement, by [DMMN24, Theorem 3] we have that
@@ Appendix A Jost solutions and intertwining operators
$K_{\kappa}\in\mathcal{C}^{\infty}(\mathbb{R}^{2}_{+}\setminus D),\qquad D=\{(x,t)\in\mathbb{R}^{2}_{+}:0<t\leq x\},$
@@
(A.4) $|K_{\kappa}(x,t)|\leq\frac{|\kappa|e^{\frac{|\kappa|}{4}}}{4}e^{-(x+t)},\qquad\text{for all }(x,t)\in\mathbb{R}^{2}_{+}\setminus D.$
Let $G_{\kappa}\in\mathcal{C}^{\infty}(D)$ be given by
@@
Moreover, $\mathcal{T}_{H}$ maps $L^{1}_{\mathrm{loc}}(\mathbb{R}_{+})$ into $\mathcal{C}(\mathbb{R}_{+})$, and $\mathcal{C}^{m}(\mathbb{R}_{+})$ into $\mathcal{C}^{m+1}(\mathbb{R}_{+})$ for every $m\in\mathbb{N}_{0}$. If, in addition, there exist $C_{H}>0$ and $g\in L^{1}_{1/2}(\mathbb{R}_{+})$ such that
@@
$G_{\kappa}^{n}(x,t)=\int_{0}^{x}\int_{0}^{x_{1}}\dots\int_{0}^{x_{n-1}}G_{\kappa}(x,x_{1})G_{\kappa}(x_{1},x_{2})\dots G_{\kappa}(x_{n-1},t)\,\mathrm{d}x_{n-1}\dots\mathrm{d}x_{2}\mathrm{d}x_{1}.$
@@
    For every $z_{0}\geq 0$, $\mathcal{T}_{\kappa}:L^{1}_{z_{0}}(\mathbb{R}_{+})\longrightarrow L^{1}_{z_{0}}(\mathbb{R}_{+})$ is bounded with norm bounded by $|\kappa|C_{\kappa}$.
@@
[…]
@@
Estimate (A.13) is a direct consequence of this. Property (ii) follows from (A.18) and Lemma A.1. To prove (iii) denote $M_{\kappa}g=\tau_{k}\star g$, we show that the Neumann series
(A.19) $\sum_{n=0}^{\infty}\mathcal{M}_{\kappa}^{n}$
@@ Appendix B Spherical harmonics and projectors
$N_{\ell,d}:=\dim\mathfrak{H}_{\ell}^{d}=\binom{\ell+d-1}{\ell}-\binom{\ell+d-3}{\ell-2}=\frac{(2\ell+d-2)(\ell+d-3)!}{\ell!(d-2)!}.$
Spherical harmonics of different degrees are orthogonal in $L^{2}(\mathbb{S}^{d-1})$. In fact
@@
$-\Delta_{\mathbb{S}^{d-1}}|_{\mathfrak{H}_{\ell}^{d}}=\ell(\ell+d-1)\operatorname{\mathrm{Id}}_{\mathfrak{H}_{\ell}^{d}},\qquad\forall\ell\in\mathbb{N}_{0}.$
@@
where $P_{\ell,d}$ are the generalized Legendre polynomials given by
$P_{\ell,d}(t)=(-1)^{\ell}R_{\ell}(1-t^{2})^{-\frac{d-3}{2}}\left(\dfrac{\mathrm{d}}{\mathrm{d}t}\right)^{\ell}(1-t^{2})^{\ell+\frac{d-3}{2}}=\frac{\ell+\nu_{d}}{\nu_{d}N_{\ell,d}}C_{\ell}^{\nu_{d}}(t),$
with $C_{\ell}^{\nu_{d}}$ a Gegenbauer polynomial of degree $\ell$ and
@@
[…]
@@
    $\int_{\mathbb{S}^{d-1}}P_{\ell,d}\left(\widehat{x}\cdot y\right)f\left(y\cdot\widehat{z}\right)\mathrm{d}y$ $=\left\lvert\mathbb{S}^{d-2}\right\rvert P_{\ell,d}\left(\widehat{x}\cdot\widehat{z}\right)\int_{-1}^{1}P_{\ell,d}(t)f(t)(1-t^{2})^{\frac{d-3}{2}}\mathrm{d}t$
    (B.4) $=\left\lvert\mathbb{S}^{d-2}\right\rvert R_{\ell}P_{\ell,d}\left(\widehat{x}\cdot\widehat{z}\right)\int_{-1}^{1}\partial_{t}^{\ell}f(t)(1-t^{2})^{\ell+\frac{d-3}{2}}\mathrm{d}t.$
@@
$=\frac{\left\lvert\mathbb{S}^{d-2}\right\rvert N_{\ell,d}R_{\ell}}{\left\lvert\mathbb{S}^{d-1}\right\rvert}P_{\ell,d}\left(x\cdot\widehat{\zeta}\right)\int_{-1}^{1}g^{(\ell)}(t)(1-t^{2})^{\ell+\frac{d-3}{2}}\mathrm{d}t$
@@
$\mathcal{P}_{\ell,d}e_{\zeta}(x)=\Gamma(d/2)N_{\ell,d}\left(\frac{2}{\sqrt{\zeta\cdot\zeta}}\right)^{\nu_{d}}I_{\ell+\nu_{d}}\left(\sqrt{\zeta\cdot\zeta}\right)P_{\ell,d}\left(x\cdot\widehat{\zeta}\right).$
@@
$\int_{\mathbb{S}^{d-1}}P_{\ell,d}\left(y\cdot\widehat{\zeta_{1}}\right)P_{\ell,d}\left(y\cdot\widehat{\zeta_{2}}\right)\mathrm{d}y$ $=\left\lvert\mathbb{S}^{d-2}\right\rvert P_{\ell,d}\left(\widehat{\zeta_{1}}\cdot\widehat{\zeta_{2}}\right)\int_{-1}^{1}P_{\ell}(t)^{2}(1-t^{2})^{\frac{d-3}{2}}\mathrm{d}t$

### v2

@@ 4.3. From the $A_{\kappa}$-amplitude to the Born approximation
    $q^{s}_{\kappa}-q\in\mathcal{C}(\overline{\mathbb{B}^{d}}\setminus\{0\})$ and $q^{s}_{\kappa}-q=0$ in $\partial\mathbb{B}^{d}$. If $q\in\mathcal{C}^{m}(\overline{\mathbb{B}^{d}}\setminus\{0\})$ for some $m\in\mathbb{N}_{0}$ then $q^{s}_{\kappa}-q$ belongs to $\mathcal{C}^{m+1}(\overline{\mathbb{B}^{d}}\setminus\{0\})$.
@@
(4.21) $\int_{\mathbb{R}_{+}}|A_{Q,\kappa}(t)|e^{-2st}\,\mathrm{d}t=\frac{1}{\left\lvert\mathbb{S}^{d-1}\right\rvert}\int_{\mathbb{B}^{d}}|q^{s}_{\kappa}(x)||x|^{2(s-\nu_{d})}\,\mathrm{d}x.$
@@
$\lambda_{\ell}[q,\kappa]-\lambda_{\ell}[0,\kappa]=-m_{Q}\left(-(\ell+\nu_{d})^{2}\right)+m_{Q_{\kappa}}\left(-(\ell+\nu_{d})^{2}\right)\\ =-\int_{0}^{\infty}A_{Q,\kappa}(t)u_{\ell+\nu_{d}}(t,\kappa)^{2}\,\mathrm{d}t=-\int_{0}^{\infty}A_{Q,\kappa}(t)e^{-2\nu_{d}t}\frac{\varphi_{\ell}(\sqrt{\kappa}e^{-t})^{2}}{\varphi_{\ell}(\sqrt{\kappa})^{2}}\,\mathrm{d}t.$
@@
which proves (4.19) when $\kappa\notin\operatorname{Spec}_{H^{1}_{0}}(-\Delta)$. To see that the result holds only under the assumption that $\kappa$ is not a Dirichlet eigenvalue of $\Delta+q$ notice that both sides of the above equations can be extended continuously to $\operatorname{Spec}_{H^{1}_{0}}(-\Delta)$ by (1.9) and Lemma 4.2.
Property (iii) follows from Theorem 4.4 (i) and (iii): $A_{Q,\kappa}-Q$ is continuous on $\mathbb{R}_{+}$ and can be extended continuously by zero to $0,\infty)$. Then $q^{s}_{\kappa}-q$ is continuous on $\mathbb{B}^{d}\setminus\{0\}$, and can be continuously extended by zero to $\overline{\mathbb{B}^{d}}\setminus\{0\}$.
Part (iv) follows from estimate (4.15) by (4.17) and a change of variables. To see this, recall that $Q\in L^{1}(\mathbb{R}_{+})$ when $p>d/2$ (see Lemma 4.1), then $\alpha_{Q}(t)\leq\left\lVert Q\right\rVert_{L^{1}(\mathbb{R}_{+})}$ for all $t\in\mathbb{R}$. Hence, there exists $\beta_{q,\kappa}>0$ such that $|x|^{-(2\alpha_{Q}(-\log|x|)+|\kappa|)}\leq|x|^{-\beta_{q,\kappa}}$.
@@ 5.1. Proof of 1 and 2
Therefore, by Theorem 3.1 we must have that
@@
which implies that $q^{\mathrm{B}}_{\kappa}\in\mathcal{B}_{d}$ (here we are using the identification of elements of $\mathcal{B}_{d}$ extended by zero, with elements of $\mathcal{E}^{\prime}_{\mathrm{rad}}(\mathbb{R}^{d})$). Finally, again by Theorem 3.1, $q^{\mathrm{B}}_{\kappa}$ satisfies (1.15). Note that once the existence of the Born approximation at energy $\kappa$ has been established, 2 follows from Proposition 4.6 and (5.3).
The proof of 1 will be concluded once we prove identity (1.17). The conditions on $\kappa$ and $q$ in the statement guarantee that the operator $\Lambda_{\kappa,q}-\Lambda_{\kappa,0}$ is well defined. Recall that $\mathcal{P}_{\ell,d}$ denotes the orthogonal projection of $L^{2}(\mathbb{S}^{d-1})$ onto $\mathfrak{H}_{\ell}^{d}$, and that $\mathcal{P}_{\ell,d}\overline{f}=\overline{\mathcal{P}_{\ell,d}f}$. Therefore we have that
@@
5.2. Proof of 3
Recall $U_{b}:=\{x\in\mathbb{B}^{d}:b<|x|<1\}$. By (4.18), (4.17) and (5.3) we have
@@ 5.2. Proof of 3
For the stability notice that if $\max_{j=1,2}\left\lVert q_{j}\right\rVert_{L^{p}(U_{b})}<K$, then one can choose a constant $M>0$ (dependent on $\kappa$) such that (4.16). Applying Theorem 4.5(ii), and letting $a=-\log b$, we obtain
@@ 5.3. Proof of 4
where $\mathcal{R}^{\mathbb{R}^{d}}_{0}(-z^{2})$ is the inverse of $\Delta+z^{2}$ defined in $L^{2}(\mathbb{R}^{d})$. This operator in the whole space has the Green function
@@
[…]
@@
[…]
@@
where $A_{Q}-A_{Q_{\kappa}}$ belongs to $L^{1}_{s}(\mathbb{R}_{+})$. Let $\psi_{z}(\cdot;\kappa)\in L^{2}(\mathbb{R}_{+})$ denote the Jost solution of (A.1). This solution is characterized by identities (A.2) and (A.3). Using Lemma A.3 iv) we can write
@@
On the other hand, by Lemma A.2 and Lemma A.3, $\mathcal{G}_{\kappa}^{-1}-\operatorname{\mathrm{Id}}$ and $\mathcal{T}_{\kappa}-\operatorname{\mathrm{Id}}$ are both Volterra operators with a smooth kernel on $D$, and thus satisfy the conditions of Lemma A.1. Therefore $A_{Q,\kappa}-(A_{Q}-A_{Q_{\kappa}})$ belongs to $\mathcal{C}([0,\infty))$ always and to $\mathcal{C}^{m+1}([0,\infty))$ if $Q\in\mathcal{C}^{m}([0,\infty))$ with $m\in\mathbb{N}_{0}$; the analogous result for $A_{Q,\kappa}-Q$ follows immediately (recall that $Q_{\kappa}$, and hence $A_{Q_{\kappa}}$, are smooth). The fact that $A_{Q,\kappa}-Q$ can be extended continuously by zero to $0,\infty)$ follows from estimate (4.15) that we now prove.
@@
Estimate (4.15) follows adding and subtracting $Q_{\kappa}-Q$ in the LHS of (6.3) and using the estimate
(6.4) $\left|A_{Q_{1}}(t)-Q_{1}(t)-\left(A_{Q_{2}}(t)-Q_{2}(t)\right)\right|\\ \leq(\alpha_{Q_{1}}(t)+\alpha_{Q_{2}}(t))e^{t(\alpha_{Q_{1}}(t)+\alpha_{Q_{2}}(t))}\int_{0}^{t}|Q_{1}(s)-Q_{2}(s)|\,\mathrm{d}s,$
twice, which can be found in [DMMN24, Section 4.1] (recall that $\alpha_{Q}$ was defined in (4.14)).
It remains to prove property (iv). If $Q$ is positive, the operator $H_{Q}=(-\partial_{t}^{2}+Q(t))$ has no negative eigenvalues nor a zero resonance, as remarked in [RS00]. Then it follows from [RS00, Theorem 3] that $A_{Q}\in L^{1}(\mathbb{R}_{+})$. Also, we have $A_{Q_{\kappa}}\in L^{1}(\mathbb{R}_{+})$ if $\kappa<(j_{0,1})^{2}$ by Lemma 4.3, since $z_{\kappa}$ can be taken to be 0. It follows that $A_{Q,\kappa}\in L^{1}(\mathbb{R}_{+})$ by (6.1), since $\mathcal{G}_{\kappa}^{-1}\mathcal{T}_{\kappa}$ is bounded on $L^{1}(\mathbb{R}_{+})$ by Lemmas A.2 and A.3. This finishes the proof of the Theorem. ∎
@@
The first statement was proved by Simon in [Sim99, Theorem 1.5] for the case $\kappa=0$. The case $\kappa\neq 0$ follows from this and (6.1) by the local injectivity of the operators $\mathcal{G}_{\kappa}^{-1}\mathcal{T}_{\kappa}$, which satisfy that $\mathcal{G}_{\kappa}^{-1}\mathcal{T}_{\kappa}F|_{(0,a)}$ vanishes iff $F|_{(0,a)}$ vanishes due to $\mathcal{G}_{\kappa}$, $\mathcal{G}_{\kappa}^{-1}$, $\mathcal{T}_{\kappa}$, $\mathcal{T}_{\kappa}^{-1}$ being Volterra operators (see Appendix A).
We now prove the stability estimate. Under the assumptions for $Q_{j}$ $j=1,2$ in the statement, by [DMMN24, Theorem 5.1] we have that
@@ Appendix A Jost solutions and intertwining operators
$K_{\kappa}\in\mathcal{C}^{\infty}(\{(x,t)\in\mathbb{R}^{2}:0\leq x\leq t\})$
@@
(A.4) $|K_{\kappa}(x,t)|\leq\frac{|\kappa|e^{\frac{|\kappa|}{4}}}{4}e^{-(x+t)},\qquad\text{for all }0\leq x\leq t.$
Set $D:=\{(x,t)\in\mathbb{R}^{2}:0\leq t\leq x\}$, let $G_{\kappa}\in\mathcal{C}^{\infty}(D)$ be given by
@@
Moreover, $\mathcal{T}_{H}$ maps $L^{1}_{\mathrm{loc}}(\mathbb{R}_{+})$ into $\mathcal{C}([0,\infty))$, and $\mathcal{C}^{m}([0,\infty))$ into $\mathcal{C}^{m+1}([0,\infty))$ for every $m\in\mathbb{N}_{0}$. If, in addition, there exist $C_{H}>0$ and $g\in L^{1}_{1/2}(\mathbb{R}_{+})$ such that
@@
$G_{\kappa}^{n}(x,t)=\int_{0}^{x}\int_{0}^{x_{1}}\dots\int_{0}^{x_{n-1}}G_{\kappa}(x,x_{1})G_{\kappa}(x_{1},x_{2})\dots G_{\kappa}(x_{n-1},t)\,\mathrm{d}x_{n-1}\dots\mathrm{d}x_{2}\mathrm{d}x_{1},$
@@
    For every $z_{0}\geq 0$, $(\mathcal{T}_{\kappa}-\operatorname{\mathrm{Id}}):L^{1}_{z_{0}}(\mathbb{R}_{+})\longrightarrow L^{1}_{z_{0}}(\mathbb{R}_{+})$ is bounded with norm bounded by $|\kappa|C_{\kappa}$.
@@
$|r_{\kappa}\star r_{\kappa}(x)|\leq\int_{0}^{x}|r_{\kappa}(x-s)||r_{\kappa}(s)|\,\mathrm{d}s\\ \leq\kappa^{2}\frac{e^{\frac{|\kappa|}{2}}}{4}\int_{0}^{x}e^{-2x}\,\mathrm{d}s=\kappa^{2}\frac{e^{\frac{|\kappa|}{2}}}{4}xe^{-2x}.$
@@
Estimate (A.14) is a direct consequence of this. Property (ii) follows from (A.18) and Lemma A.1. To prove (iii) denote $\mathcal{M}_{\kappa}g=\tau_{k}\star g$, we show that the Neumann series
(A.19) $\sum_{n=0}^{\infty}(-1)^{n}\mathcal{M}_{\kappa}^{n}$
@@ Appendix B Spherical harmonics and projectors
$N_{\ell,d}:=\dim\mathfrak{H}_{\ell}^{d}=\binom{\ell+d-1}{\ell}-\binom{\ell+d-3}{\ell-2}=\frac{(\ell+d-2)!+\ell(\ell+d-3)!}{\ell!(d-2)!}.$
Spherical harmonics of different degrees are orthogonal in $L^{2}(\mathbb{S}^{d-1})$. In fact,
@@
$\Delta_{\mathbb{S}^{d-1}}|_{\mathfrak{H}_{\ell}^{d}}=\ell(\ell+d-2)\operatorname{\mathrm{Id}}_{\mathfrak{H}_{\ell}^{d}},\qquad\forall\ell\in\mathbb{N}_{0}.$
@@
where $P_{\ell,d}$ are the generalized Legendre polynomials
$P_{\ell,d}(t)=(-1)^{\ell}R_{\ell}(1-t^{2})^{-(\nu_{d}-\frac{1}{2})}\left(\dfrac{\mathrm{d}}{\mathrm{d}t}\right)^{\ell}(1-t^{2})^{\ell+\nu_{d}-\frac{1}{2}},$
and
@@
(B.2) $\frac{|\mathbb{S}^{d-2}|}{|\mathbb{S}^{d-1}|}\int_{-1}^{1}P_{m,d}(t)P_{n,d}(t)(1-t^{2})^{\nu_{d}-\frac{1}{2}}\mathrm{d}t=\frac{1}{N_{m,d}}\delta_{m,n}.$
@@
    $\int_{\mathbb{S}^{d-1}}P_{\ell,d}\left(\widehat{x}\cdot y\right)f\left(y\cdot\widehat{z}\right)\mathrm{d}y$ $=\left\lvert\mathbb{S}^{d-2}\right\rvert P_{\ell,d}\left(\widehat{x}\cdot\widehat{z}\right)\int_{-1}^{1}P_{\ell,d}(t)f(t)(1-t^{2})^{\nu_{d}-\frac{1}{2}}\mathrm{d}t$
    (B.4) $=\left\lvert\mathbb{S}^{d-2}\right\rvert R_{\ell}P_{\ell,d}\left(\widehat{x}\cdot\widehat{z}\right)\int_{-1}^{1}\partial_{t}^{\ell}f(t)(1-t^{2})^{\ell+\nu_{d}-\frac{1}{2}}\mathrm{d}t.$
@@
$=\frac{\left\lvert\mathbb{S}^{d-2}\right\rvert N_{\ell,d}R_{\ell}}{\left\lvert\mathbb{S}^{d-1}\right\rvert}P_{\ell,d}\left(x\cdot\widehat{\zeta}\right)\int_{-1}^{1}g^{(\ell)}(t)(1-t^{2})^{\ell+\nu_{d}-\frac{1}{2}}\mathrm{d}t$
@@
[…]
@@
$\int_{\mathbb{S}^{d-1}}P_{\ell,d}\left(y\cdot\widehat{\zeta_{1}}\right)P_{\ell,d}\left(y\cdot\widehat{\zeta_{2}}\right)\mathrm{d}y$ $=\left\lvert\mathbb{S}^{d-2}\right\rvert P_{\ell,d}\left(\widehat{\zeta_{1}}\cdot\widehat{\zeta_{2}}\right)\int_{-1}^{1}P_{\ell}(t)^{2}(1-t^{2})^{\nu_{d}-\frac{1}{2}}\mathrm{d}t$